infinite wire is bent into a simicircule shown on the right with Pwhen the current in the wire is i, witt be mag 41 - . adius ρ eid atPt P2
Answer #1

We need to find the magnetic moment which is given as,

M=NIA

Where, N= Number of turns.

I= Current in the coil.

A= Area of the coil,which is a circle.

Now try to understand,

When the wire of length L is bent in the form of the circle, the perimeter of the circle will be L.

So, 2πr=L

⇒r=L/2π

So, you can find the area now which is,

A=πr2=π×L2 /4π2=L/4π

Here N=1

So magnetic moment has a magnitude of,

Direction :-

Suppose the coil is on the plane of the paper.

If the direction of current is clockwise then according to right hand thumb rule magnetic moment will be

We need to find the magnetic moment which is given as, M-NIA where, N= Number of turns. Current in the coil. A= Area of the coil,which is a circle Now try to understand, When the wire of length L is bent in the form of the circle, the perimeter of the circle will be L So, 2Ttr-L So you can find the area now which is Here N=1 So magnetic moment has a magnitude of, M-NIL 2/4T Direction: Suppose the coil is on the plane of the paper. · lf the direction of current is clockwise then according to right hand thumb rule magnetic moment will be inward normal to the plane of paper(means x magnetic moment) If the direction of current is anticlockwise then the direction of magnetic moment is outward normal to the plane of paper(means .magnetic moment)